OpenStudy (anonymous):

Decide whether each of the following is linear or not linear. In each case, explain your answer. a. 3x+4y+2=0 x | y 2 1 4 5 6 15 8 30

OpenStudy (anonymous):

c. You have a secret that you tell to one person. Every hour, each of the people that know the secret tells one person. The number of people who know is N, and t is the number of hours since you told the first person. Is N a linear function of t?

OpenStudy (anonymous):

sorry forgot to put the B in front of the second one lol

jagr2713 (jagr2713):

@freckles @OOOPS

OpenStudy (anonymous):

not linear

OpenStudy (anonymous):

why ? @OOOPS

OpenStudy (anonymous):

when you draw this relation u will got a straight line

OpenStudy (anonymous):

there are three questions here a b c, i think you only got the first one if im not mistaken

OpenStudy (freckles):

I'm not really sure what's going on here. a. is 3x+4y+2=0 b. x|y c. is that last one you put that wasn't included in the original? what is that other stuff 2 1 4 5 6 15 8 30 2 doesn't divide 1 4 doesn't divide 5 and so on so i know these aren't examples of x|y

OpenStudy (anonymous):

yes b is the x y chart a is that c is the last response after the first one lol

OpenStudy (anonymous):

I thought those are the points in 3x+4y+2=0, that 's why I said it is not linear.:)

OpenStudy (anonymous):

the x y chart is the 2 1 and 4 5

OpenStudy (anonymous):

the left is x the right is y XD

OpenStudy (freckles):

a) well ax+by+c=0 is a linear equation form b) {(x,y)|(2,1),(4,5),(6,15),(8,30)} If this is a linear function then \[\frac{5-1}{4-2}=\frac{15-1}{6-2}=\frac{30-1}{8-2}=\frac{15-5}{6-4}=\frac{30-5}{8-4}=\frac{30-15}{8-6}\] or you could find the equation of the line involving the first 2 pairs only then plug in the other order pairs to see they belong on that line example: \[\frac{5-1}{4-2}=\frac{4}{2}=2\] \[y-1=2(x-2) \text{ point slope form }\] then take (4,5) and see if it is on that line

OpenStudy (anonymous):

thanks for trying to help freckles, read pm :)

OpenStudy (freckles):

c. You have a secret that you tell to one person. Every hour, each of the people that know the secret tells one person. The number of people who know is N, and t is the number of hours since you told the first person. Is N a linear function of t? for the last one... So at t_0 hour we have N people know at t_1 we would have 2N at t_2 we would have 2(2N)=4N the amount is just doubling... So this should be a hint for the last one.

OpenStudy (anonymous):

thank you